# Non-intersecting Squared Bessel Process: Spectral Moments and Dynamical Entanglement Entropy

### v1

@@ 1 Introduction
Quantifying entanglement is a central problem in quantum information theory, with direct implications for quantum communication, computation, and many-body physics. For bipartite quantum systems in random pure states, the reduced density matrix of a subsystem exhibits highly nontrivial spectral fluctuations, characterized by strong eigenvalue correlations, which are naturally described within random matrix theory [57, 25, 77, 54, 55]. In Page’s seminal work [60], the eigenvalues of reduced density matrices follow the Hilbert-Schmidt ensemble (fixed-trace Wishart-Laguerre ensemble), where the degree of entanglement is estimated via cumulants of entanglement metrics such as entanglement entropy [60, 30, 61, 65, 67, 68, 34, 38], quantum purity [51, 76, 31], and entanglement capacity [58, 71]. Similar statistical formulations for other random state ensembles have been considered in the literature, notably the Bures-Hall ensemble [63, 59, 7, 8, 6, 9, 24, 62, 70, 69, 53, 72] and fermionic Gaussian ensemble [2, 3, 35, 36, 37]. Despite substantial progress in those statistical settings, such results are inadequate for realistic quantum systems, where entanglement is inherently non-stationary. Studies of entanglement estimation under real-world scenario of time-dependent stochastic models remain relatively limited. Time-dependent dynamical random matrix ensembles have, however, been extensively studied including most notably the Dyson Brownian motion and its extensions [22, 4, 43, 44, 45, 48, 25, 46]. Dynamical random matrix framework provides a natural and well-defined mechanism that is well suited for introducing time-dependent deformations into statistical quantum state models to investigate the evolution of quantum entanglement. Understanding this behavior helps identify statistical features of entanglement that are relevant for realistic quantum applications, where entanglement evolves over time during state preparation and algorithm execution.
@@
Technically, rather than relying on direct, case-by-case computations [30, 61, 63, 65, 67, 62, 68, 70, 69, 34, 53, 35, 36, 71, 37, 38, 72], these results follow from new recurrence relations for spectral moments derived in this work. Spectral moments is a fundamental object in random matrix theory [40, 39, 50, 21, 74, 32, 16, 26, 27, 18, 75, 10, 11, 12, 1, 13, 28], encoding global information about eigenvalue distributions and underpinning many classical results such as Wigner’s semicircle law [73], the Marchenko–Pastur law for sample covariance matrices [56], genus expansions and map enumeration arising from exact trace moment formulas [40], free convolution limits and asymptotic freeness in free probability [66], central limit theorems for linear eigenvalue statistics [41], and strong convergence results for random matrices [39, 50]. Spectral moments often satisfy exact recurrence relations, see for example, the prototypical cases in [40, 39, 50]. These recursions provide a powerful tool for both finite-size computations and asymptotic analysis of random matrix ensembles. While existing works have mostly focused on spectral moments of integer order, we extend this framework to real orders, naturally leading to logarithmic linear statistics relevant to entanglement entropy. This approach builds on the framework recently developed in [38], where spectral moments serve as fundamental building blocks for the systematic derivation of higher-order cumulants of entanglement entropy in the fixed-trace Wishart–Laguerre ensemble.
@@ 1.1 Non-intersecting squared Bessel process
We introduce here the formulation that leads to the non-intersecting squared Bessel process [14, 15, 43, 44, 45, 29, 48, 25, 46, 19, 20, 49, 47, 52]. Let $X(t)$ be a one-dimensional diffusion with transition probability density $\widehat{P}_{t}(x,y)$. By the Karlin-McGregor theorem [42], the joint probability density of positions $(x_{1},\dots,x_{n})=(X_{1}(t),\dots,X_{n}(t))$ of $n$ independent copies $X_{j}$, $j=1,\dots,n$ with endpoints
@@ A Christoffel-Darboux formula of correlation kernel (15) is given by
where $f(x,y)$ takes the form (87) and g(x,y)=xyg(x,y)=x-y. This limit
@@ 2.2 Recurrence relations of spectral moments
$m_{k}^{\pm l}=\!\left.\mathbb{E}[R_{k}]\right\rvert_{n\to n\pm l}.$ (102)
@@ 3 Dynamical Entanglement Estimation
In the present dynamical model, the joint density (108) does not factorize into independent distributions for $r$ and $\boldsymbol{\lambda}$, preventing the exact moment conversions. To proceed, we replace the random variable $r$ by its mean value $m_{1}$, where we recall $m_{k}$ defined in (13). Under this approximation, entanglement metrics can be expressed in terms of linear statistics of the variables $x_{i}$. This approximation is motivated by the canonical–microcanonical equivalence principle familiar from statistical mechanics, where a fluctuating extensive constraint is replaced by its typical value when concentration holds [23, 64].

### v2

@@ 1 Introduction
Quantifying entanglement is a central problem in quantum information theory, with direct implications for quantum communication, computation, and many-body physics. For bipartite quantum systems in random pure states, the reduced density matrix of a subsystem exhibits highly nontrivial spectral fluctuations, characterized by strong eigenvalue correlations, which are naturally described within random matrix theory [59, 27, 56, 57]. In Page’s seminal work [62], the eigenvalues of reduced density matrices follow the Hilbert-Schmidt ensemble (fixed-trace Wishart-Laguerre ensemble), where the degree of entanglement is estimated via cumulants of entanglement metrics such as entanglement entropy [62, 32, 63, 67, 69, 70, 36, 40], quantum purity [53, 77, 33], and entanglement capacity [60, 73]. Similar statistical formulations for other random state ensembles have been considered in the literature, notably the Bures-Hall ensemble [65, 61, 7, 8, 6, 9, 26, 64, 72, 71, 55, 74] and fermionic Gaussian ensemble [2, 3, 37, 38, 39]. Despite substantial progress in those statistical settings, such results are inadequate for realistic quantum systems, where entanglement is inherently non-stationary. Studies of entanglement estimation under real-world scenario of time-dependent stochastic models remain relatively limited. Time-dependent dynamical random matrix ensembles have, however, been extensively studied including most notably the Dyson Brownian motion and its extensions [24, 4, 45, 46, 47, 50, 27, 48, 20]. Dynamical random matrix framework provides a natural and well-defined mechanism that is well suited for introducing time-dependent deformations into statistical quantum state models to investigate the evolution of quantum entanglement. Understanding this behavior helps identify statistical features of entanglement that are relevant for realistic quantum applications, where entanglement evolves over time during state preparation and algorithm execution.
@@
Technically, rather than relying on direct, case-by-case computations [32, 63, 65, 67, 69, 64, 70, 72, 71, 36, 55, 37, 38, 73, 39, 40, 74], these results follow from new recurrence relations for spectral moments derived in this work. Spectral moments is a fundamental object in random matrix theory [42, 41, 51, 52, 23, 17, 34, 17, 29, 30, 19, 76, 10, 11, 12, 1, 13, 31], encoding global information about eigenvalue distributions and underpinning many classical results such as Wigner’s semicircle law [75], the Marchenko-Pastur law for sample covariance matrices [58], genus expansions and map enumeration arising from exact trace moment formulas [42], free convolution limits and asymptotic freeness in free probability [68], central limit theorems for linear eigenvalue statistics [43], and strong convergence results for random matrices [41, 52]. Spectral moments often satisfy exact recurrence relations, see for example, the prototypical cases in [42, 41, 52]. These recursions provide a powerful tool for both finite-size computations and asymptotic analysis of random matrix ensembles. While existing works have mostly focused on spectral moments of integer order, we extend this framework to real orders, naturally leading to logarithmic linear statistics relevant to entanglement entropy. This approach builds on the framework recently developed in [40], where spectral moments serve as fundamental building blocks for the systematic derivation of higher-order cumulants of entanglement entropy in the fixed-trace Wishart–Laguerre ensemble.
@@ 1.1 Non-intersecting squared Bessel process
We introduce here the formulation that leads to the non-intersecting squared Bessel process [14, 15, 45, 46, 47, 28, 50, 27, 48, 21, 21, 49, 54]. Let $X(t)$ be a one-dimensional diffusion with transition probability density $\widehat{P}_{t}(x,y)$. By the Karlin-McGregor theorem [44], the joint probability density of positions $(x_{1},\dots,x_{n})=(X_{1}(t),\dots,X_{n}(t))$ of $n$ independent copies $X_{j}$, $j=1,\dots,n$ with endpoints
@@ A Christoffel-Darboux formula of correlation kernel (15) is given by
where $f(x,y)$ takes the form (87) and g(x,y)=xyg(x,y)=x-y. The limit
@@ 2.2 Recurrence relations of spectral moments
$m_{k}^{\pm l}=\mathbb{E}[R_{k}]\rvert_{n\to n\pm l}.$ (102)
@@ 3 Dynamical Entanglement Estimation
In the present dynamical model, the joint density (108) does not factorize into independent distributions for $r$ and $\boldsymbol{\lambda}$, preventing the exact moment conversions. To proceed, we replace the random variable $r$ by its mean value $m_{1}$, where we recall $m_{k}$ defined in (13). Under this approximation, entanglement metrics can be expressed in terms of linear statistics of the variables $x_{i}$. This approximation is motivated by the canonical-microcanonical equivalence principle familiar from statistical mechanics, where a fluctuating extensive constraint is replaced by its typical value when concentration holds [25, 66].